### v1

@@ 3.4 Comparison between AGN and Galactic black hole binaries
The comparison between AGN and GBHB PSDs is complicated by the fact that the GBHB PSDs depend on the “state” in which they operate [e.g. 26, 254]. The right panel in Fig. 28 shows the power spectrum of Cyg X-1 in its “canonical soft state”, taken from [4]. As we discussed above, this PSD is very similar to the AGN PSDs (both in shape and in amplitude). However, as [4] notice, only 35% of the observations made during the soft state of Cyg X-1 allowed a fit using a simple cutoff PL model (indicated by the solid line in this plot).
@@ 3.5 Quasi-periodic oscillations (QPOs) in AGN
QPOs are commonly observed in the X-ray light curves of Galactic X-ray binaries, both in black hole and in neutron star systems. QPOs are detected using Fourier techniques. They appear in the power spectrum as narrow peaks. By definition, their width is (at least) less than half of their centroid frequency. In GBHBs, QPOs are generally grouped into low-frequency (LF) QPOs, with centroid frequency Hz, and high-frequency (HF) QPOs, with centroid frequency Hz [e.g. 26]. LF QPOs in GBHBs have been classified into three types: A, B and C. In general, Type C QPOs are observed during the start of the outburst when GBHBs are in their hard state (but also appear in later stages), whereas Type A and Type B QPOs are only observed during the transition to the soft state [119].
@@
HF QPOs are rare in GBHBs [e.g. 39]. They appear only in high-flux observations, with a fairly specific hardness ratio [119]. According to [254], the HF QPOs appear during the “Very High” or “Steep Power law” state of GBHBs, where the PSD amplitude is (in the PSD representation). This is significantly lower than the mean PSD amplitude of AGN. Therefore, it is not surprising that QPOs are very rare in AGN as well. For example, [102] find no strong evidence for QPOs in their analysis of archival XMM-Newton light curves, except for the well-known case of RE J1034+396. The QPO in this galaxy was first reported by [97], and has been reconfirmed many times since then [e.g. 11, 121, 331, 318].
Detections of similar QPOs in other AGN [e.g. 14, 248, 333, 101] are less convincing. The main issue with these detections is the transient nature of these QPOs. They have been detected in only one XMM-Newton observation in all cases, casting some doubt on the validity of the detections. For example, a period of s in AGN (this is the typical duration of an XXM-Newton observation) will correspond to a period of s in GBHBs (if we assume a BH mass of M⊙ in AGN and of M⊙ in GBHBs). If real, the QPOs in AGN should correspond to QPOs that would appear only for a period of time not more than s in GBHBs (and then they would disappear). Such QPOs have never been observed (although, even if they existed, we would not be able to detect them in the GBHB light curves).
@@ 3.6.1 The high frequency PSD slope
As discussed above, the low-frequency PSD slope is below the PSD bending frequency in the 2–10 keV band. This result also holds at higher energies. [217] found that the low-frequency slope of AGN PSDs in the 15-190 keV band is also close to . However, this is not the case for the PSD slope at frequencies higher than . [228] were the first to notice that the slope of the PSD becomes flatter (that is, the PSD “hardens”) with increasing energy, when studying the X-ray PSD of NGC 4051 below and above 2 keV, using EXOSAT light curves. This effect has been observed in many more AGN and is similar to what has also been observed in GBHBs [e.g. 213, for Cyg X-1].
The left panel in Fig. 29 shows the distribution of the best-fit PSD slopes from [102]. The black dashed vertical lines indicate the mean PSD slope when the power spectra are well fitted by a PL model. On average, the mean slope of the 2–10 keV band PSDs is flatter than the one in the 0.2-2 keV band (middle and bottom panels in the same Figure). Interestingly, this is not the case for the sources where the BPL model fits the power spectrum well. The high-frequency slope in this case is steeper than the slope in the case of the PL models, and it does not appear to depend on energy (see the red dashed vertical lines in the same figure). It is not clear whether this result depends on the fact that, as we have already commented, it may not be easy to accurately determine the PSD properties by using XMM-light curves only, especially in narrow energy bands, when the Poisson noise level will increase. On the other hand, the analysis of high- samples by Paolillo et al. [236] suggests that the high-frequency PSD slope in the rest frame keV band is , in excellent agreement with [102]. A detailed study of the X-ray variability properties of NGC 4051 using NICER light curves showed that the slope of the high-frequency PSD is at all energies from to 3 keV. This is an order of magnitude in frequencies, and the high signal-to-noise NICER light curves allow the determination of the PSD at very high frequencies. So, it is not clear yet whether the high-frequency PSD slope depends on energy or not.
@@ 3.6.3 The PSD Amplitude
The results of the PSD analysis of many Seyferts using RXTE light curves suggest that, on average, PSD is roughly flat and equal to at frequencies below [see also 236]. [217] found that PSD is also flat and equal to at low frequencies in the 14-195 keV energy band. The PSD amplitude then does not appear to significantly depend on energy (if at all).
This result has interesting implications on the nature of the X-ray variability in AGN. Power spectra in AGN are (almost) always divided by the square of the mean of the light curve. As a result, the PSD does not have units of flux2/Hz but units of Hz-1. In this way, we can compare PSDs that are computed using light curves from different X-ray instruments. However, the PSD amplitude is affected by this normalisation. If there are spectral components that are constant in an energy band, then the PSD amplitude will be reduced. Furthermore, if the observed variations are not due to the X-ray continuum variations only (e.g. variable absorption could affect the observed variations at low energies), then the PSD amplitude will increase. The fact that the mean PSD amplitude of the X-ray bright AGN is very similar in the 2–10 and 14-195 keV bands indicates that the main driver of the observed variations over almost two decades in energy is the X-ray continuum variations, whose amplitude does not depend on energy.
@@ 3.7 The energy-dependent X-ray time-lags in AGN
Studying time delays in the time domain is not easy. The points in the sampled cross-correlation function are heavily correlated (because the light curve points are correlated themselves). Therefore, cross-correlation functions cannot easily be used to fit theoretical models (see Sect. 3.2). The average time delay can be determined by various Monte Carlo techniques (such as ICCF, for example; 224). However, the delays between X-ray variations in various energy bands are more complicated than just a constant delay that increases with energy separation. It turns out that if we decompose the light curves into sinusoids, we can show that there is a difference between the phase of these sinusoids at each frequency that increases with decreasing frequency. When divided by 2, this phase shift gives us the time delay between each of these variability components in two energy bands.
The first detection of these frequency-dependent time delays in AGN (the so-called “hard” time lags) was presented by [232]. They studied RXTE light curves of NGC 7469 and found that the delay between the variations detected in the 4–10 and 2–10 keV bands is not constant. Instead, it increases with increasing Fourier period (left panel of Fig. 31). The dashed line in this plot shows the best fit power-law model to the data (with slope of ). The delays are on the order of % of the Fourier period. These results are similar to what has also been detected in Galactic black hole and neutron star binaries. Hard X-rays delayed with respect to softer X-rays, with a time delay approximately proportional to the Fourier period, have also been detected in many other AGN since then [e.g. 306, 171, 174, 196, 12, 259, 317, 7, 154, 165, 192, 161].
The right panel of Fig. 31 shows time lags versus the Fourier period for NGC 4051 [199]. Time-lags in this source have been studied extensively. [20] found that the delays may depend on the source flux, while [243] found that the time delays between the hard and soft band variations may extend to very long periods (of the order of hundreds of days). A comparison between the graph on the right panel of Fig. 31 and the cross-correlation plot in the right panel of Fig. 30 reveals the power of studying the delays in the frequency domain instead of the time domain. The cross-correlation function indicates a constant delay between the hard- and soft-band variations, whereas the time lags reveal that the delays between the sinusoids that are responsible for the observed variations are not constant. Instead, the delays increase with increasing Fourier period, and for any given period, the time lag increases with the energy separation of the bands. As we mentioned, this behaviour can be explained by the Comptonization hypothesis, but in this case, we would expect that the delays between various energy bands should increase by the same amount for all Fourier periods, which is not what is observed [e.g. 16].
@@
This statistic is an estimate of the integral of the power spectrum from the shortest to the longest frequency sampled by the light curve. It is a biased estimator of this integral (even in the case of evenly sampled light curves), because of the red noise character of the intrinsic X-ray variations, and the bias depends on the PSD slope and the sampling pattern of the light curve. Its statistical properties have been studied by Vaughan et al. [305, also see ], in the case of evenly sampled light curves, and by [15], in all cases. In general, a single measurement of the normalised excess variance is not very useful. Even when corrected for bias, single measurements follow a very asymmetric and broad distribution of unknown variance (that is, their error is unknown), especially in the case of sparse sampling and low signal-to-noise light curves. If we can calculate excess variance using light curves from many individual sources, then ’ensemble’ estimates of can be used, as they have a known error and distributions similar to a Gaussian [15]. The same procedure can be followed with excess variance measurements using segments of long light curves of a single object. However, the individual values in this case are not independent (due to the strong correlation between the points of the light curve over very long periods), therefore it will be difficult to determine the statistical properties of the resulting ensemble estimates.
@@
[234] computed for a large sample of (nearby) AGN using XMM-newton light curves, to study the X-ray variability of AGN in various energy bands and on timescales shorter than a day. They were able to show that their results could only be explained if the PSD bending frequency depends on both the BH mass and the accretion rate as shown by McHardy et al. [193, see the right panel of Fig. 32]. They also suggested that the PSD amplitude should increase with decreasing accretion rate in AGN. Although they predicted a rather steep dependence of the PSD amplitude on , which is not supported by other observations, their results showed that analysis can provide powerful insights into the properties of AGN X-ray variability. In fact, since can be used to characterise the X-ray variability of large samples of sources, the results of the studies can impose constraints on models for fundamental relations in AGN. For instance, [98] linked the demographics of AGN to their ensemble X-ray variability properties in order to investigate the black hole mass versus stellar mass relation in these objects (also see 280 for a similar application using PSD analysis).
@@
Given the tight correlation between and MBH, it was soon suggested that could be used to estimate the BH mass in AGN [e.g. 211, 207, 206]. [234] provided BH mass estimates for tens of AGN based on their measurements, and [332] provided a calibration for the correlation between and MBH in AGN. [13] critically assessed as a tool to measure MBH in AGN. They showed that it is possible to accurately measure MBH using measurements in the 3–10 and the 10–20 keV bands; however, the signal-to-noise ratio and the duration of the light curves should be larger than and ksec, respectively.
@@
All the works mentioned above used light curves that are shorter than day. Excess-variance measurements have perhaps played an even more important role in the study of the X-ray variability of high- AGN on longer timescales. In most cases, such objects are observed only a few times, in a highly irregular pattern. Consequently, the resulting light curves are not appropriate for the application of Fourier techniques. Therefore, has been widely used in the study of their X-ray variability properties. The lack of BH mass estimates was another obstacle in investigating the correlations between the amplitude of the X-ray variability and the physical parameters of AGN (i.e. MBH and ). As a result, most studies originally tried to investigate the correlation between variability amplitude and source luminosity, as well as redshift [e.g. 10, 170, 244, 220, 176, 320, 260, 319, 235, 330], confirming the anticorrelation with luminosity observed at lower redshift. This finding was also confirmed by studies based on SF analysis [310]. Almaini et al. [10], Manners et al. [170], Paolillo et al. [244] also observed a tentative evolution of X-ray variability with redshift, which was however disproved by later studies using better data and more complete samples.
@@ 3.9 Stationarity
It is well-known that the X-ray variability process in AGN is not strictly stationary because the variance of the lightcurves increases linearly with source flux. This “rms-flux” relation was first noticed by [299], and it is well established in AGN as well as in GBHBs [e.g. 116]. This is an important property and can place constraints on various models that have been proposed to explain the observed variations. For example, [303] showed that if the rms-flux relation holds for variations on all timescales, then the variability process in AGN should be multiplicative in nature: variations should be coupled together on all timescales. Therefore, the observed variability cannot be due to a shot-noise process or the result from completely independent variations in separate emitting regions.
@@ 3.10 Constrains on theoretical models
According to the model, the power spectra should extend with a slope of over a frequency range determined by up to 1/ (where the PSD break should appear). However, should be very large in order to explain power spectra such as the NGC 4051 PSD (left panel in Fig. 26). This power spectrum extends with a slope of over more than five orders in frequency down to Hz. See also the high-quality X-ray PSD of NGC 4395 in [32], which extends with a slope of below over four and a half orders of magnitudes in frequency. In order to explain this, the hot corona must extend to very large radii. For example, [251] fitted the Cyg X-1 PSD in its soft state with a fluctuations propagation model and found that the corona (which is located on top and below the disc) should have a very large radius of . This is clearly much larger than the size of the X-ray corona, as measured by monitoring observations of lensed quasars (see references at the beginning of Sect. 3). Apart from the difference between the model and the observed corona size, it is also difficult to understand how such a corona can be formed at such large radii from the BH, whether it will be stable or how it may evolve over time. Currently, there are no physical models that could be used to determine the accretion rate of the corona, its temperature, and the density radial distribution.
Note that the requirement of an extended corona does not come from the need to explain the observed long-term variations since these variations COULD be produced by accretion rate fluctuations which could propagate in the accretion disc and then within the corona as well). However, an accreting X-ray corona, with a large radius, is necessary to explain the long-term, continuum (hard) X-ray time-lags in AGN, especially if they really extend to very low frequencies [e.g. 243]. We also stress the similarity between the AGN PSDs and that of Cyg X-1 in its (own) soft state. In this state, a strong black-body-like component appears at low energies in the energy spectrum of the source. It is generally accepted that this component is due to thermal emission from the accretion disc, which extends down to the innermost stable circular orbit. If the similarity of the timing properties implies similar physical conditions as well, then it is possible that the accretion disc in the AGN also extends to its innermost radius, leaving not much space for a radially extended corona.

### v3

@@ 3.4 Comparison between AGN and Galactic black hole binaries
The comparison between AGN and GBHB PSDs is complicated by the fact that the GBHB PSDs depend on the “state” in which they operate [e.g. 27, 256]. The right panel in Fig. 28 shows the power spectrum of Cyg X-1 in its “canonical soft state”, taken from [4]. As we discussed above, this PSD is very similar to the AGN PSDs (both in shape and in amplitude). However, as [4] notice, only 35% of the observations made during the soft state of Cyg X-1 allowed a fit using a simple cutoff PL model (indicated by the solid line in this plot).
@@ 3.5 Quasi-periodic oscillations (QPOs) in AGN
QPOs are commonly observed in the X-ray light curves of Galactic X-ray binaries, both in black hole and in neutron star systems. QPOs are detected using Fourier techniques. They appear in the power spectrum as narrow peaks. By definition, their width is (at least) less than half of their centroid frequency. In GBHBs, QPOs are generally grouped into low-frequency (LF) QPOs, with centroid frequency Hz, and high-frequency (HF) QPOs, with centroid frequency Hz [e.g. 27]. LF QPOs in GBHBs have been classified into three types: A, B and C. In general, Type C QPOs are observed during the start of the outburst when GBHBs are in their hard state (but also appear in later stages), whereas Type A and Type B QPOs are only observed during the transition to the soft state [121].
@@
HF QPOs are rare in GBHBs [e.g. 40]. They appear only in high-flux observations, with a fairly specific hardness ratio [121]. According to [256], the HF QPOs appear during the “Very High” or “Steep Power law” state of GBHBs, where the PSD amplitude is (in the PSD representation). This is significantly lower than the mean PSD amplitude of AGN. Therefore, it is not surprising that QPOs are very rare in AGN as well. For example, [104] find no strong evidence for QPOs in their analysis of archival XMM-Newton light curves, except for the well-known case of RE J1034+396. The QPO in this galaxy was first reported by [99], and has been reconfirmed many times since then [e.g. 12, 123, 333, 320].
Detections of similar QPOs in other AGN [e.g. 15, 250, 335, 103] are less convincing. The main issue with these detections is the transient nature of these QPOs. They have been detected in only one XMM-Newton observation in all cases, casting some doubt on the validity of the detections. For example, a period of s in AGN (this is the typical duration of an XXM-Newton observation) will correspond to a period of s in GBHBs (if we assume a BH mass of M⊙ in AGN and of M⊙ in GBHBs). If real, the QPOs in AGN should correspond to QPOs that would appear only for a period of time not more than s in GBHBs (and then they would disappear). Such QPOs have never been observed (although, even if they existed, we would not be able to detect them in the GBHB light curves).
@@ 3.6.1 The high frequency PSD slope
As discussed above, the low-frequency PSD slope is below the PSD bending frequency in the 2–10 keV band. This result also holds at higher energies. [219] found that the low-frequency slope of AGN PSDs in the 15-190 keV band is also close to . However, this is not the case for the PSD slope at frequencies higher than . [230] were the first to notice that the slope of the PSD becomes flatter (that is, the PSD “hardens”) with increasing energy, when studying the X-ray PSD of NGC 4051 below and above 2 keV, using EXOSAT light curves. This effect has been observed in many more AGN and is similar to what has also been observed in GBHBs [e.g. 215, for Cyg X-1].
The left panel in Fig. 29 shows the distribution of the best-fit PSD slopes from [104]. The black dashed vertical lines indicate the mean PSD slope when the power spectra are well fitted by a PL model. On average, the mean slope of the 2–10 keV band PSDs is flatter than the one in the 0.2-2 keV band (middle and bottom panels in the same Figure). Interestingly, this is not the case for the sources where the BPL model fits the power spectrum well. The high-frequency slope in this case is steeper than the slope in the case of the PL models, and it does not appear to depend on energy (see the red dashed vertical lines in the same figure). It is not clear whether this result depends on the fact that, as we have already commented, it may not be easy to accurately determine the PSD properties by using XMM light curves only, especially in narrow energy bands, when the Poisson noise level will increase. On the other hand, the analysis of high- samples by Paolillo et al. [238] suggests that the high-frequency PSD slope in the rest frame keV band is , in excellent agreement with [104]. A detailed study of the X-ray variability properties of NGC 4051 using NICER light curves showed that the slope of the high-frequency PSD is at all energies from to 3 keV. This is an order of magnitude in frequencies, and the high signal-to-noise NICER light curves allow the determination of the PSD at very high frequencies. So, it is not clear yet whether the high-frequency PSD slope depends on energy or not.
@@ 3.6.3 The PSD Amplitude
The results of the PSD analysis of many Seyferts using RXTE light curves suggest that, on average, PSD is roughly flat and equal to at frequencies below [see also 238]. [219] found that PSD is also flat and equal to at low frequencies in the 14-195 keV energy band. The PSD amplitude does not appear to significantly depend on energy (if at all).
This result has interesting implications on the nature of the X-ray variability in AGN. Power spectra in AGN are (almost) always divided by the square of the mean of the light curve. As a result, the PSD does not have units of flux2/Hz but units of Hz-1. In this way, we can compare PSDs that are computed using light curves from different X-ray instruments. However, the PSD amplitude is affected by this normalisation. If there are spectral components that are constant in an energy band, then the PSD amplitude will be reduced. Furthermore, if the observed variations are not due to the X-ray continuum variations only (e.g. variable absorption could affect the observed variations at low energies), then the PSD amplitude will increase. The fact that the mean PSD amplitude of the X-ray bright AGN is very similar in the 2–10 and 14-195 keV bands indicates that the main driver of the observed variations over almost two decades in energy is the intrinsic X-ray continuum variations, whose amplitude does not depend on energy.
@@ 3.7 The energy-dependent X-ray time-lags in AGN
Studying delays in the time domain is not easy. The points in the sampled cross-correlation function are heavily correlated (because the light curve points are correlated themselves). Therefore, cross-correlation functions cannot easily be used to fit theoretical models (see Sect. 3.2). The average time delay between light curves in different energy bands can be determined by various Monte Carlo techniques (such as ICCF, for example; 226). However, the delays between X-ray variations in various energy bands are more complicated than just a constant delay that increases with energy separation. It turns out that if we decompose the light curves into sinusoids, we can show that there is a difference between the phase of these sinusoids at each frequency that increases with decreasing frequency. When divided by 2, this phase shift gives us the time delay between each of these variability components in two energy bands.
The first detection of these frequency-dependent time delays in AGN (the so-called “hard” time lags) was presented by [234]. They studied RXTE light curves of NGC 7469 and found that the delay between the variations detected in the 4–10 and 2–10 keV bands is not constant. Instead, it increases with increasing Fourier period (left panel of Fig. 31). The dashed line in this plot shows the best fit power-law model to the data (with slope of ). The delays are on the order of % of the Fourier period. These results are similar to what has also been detected in Galactic black hole and neutron star binaries. Hard X-rays delayed with respect to softer X-rays, with a time delay approximately proportional to the Fourier period, have also been detected in many other AGN since then [e.g. 308, 173, 176, 198, 13, 261, 319, 8, 156, 167, 194, 163].
The right panel of Fig. 31 shows time lags versus the Fourier period for NGC 4051 [201]. Time-lags in this source have been studied extensively. [21] found that the delays may depend on the source flux, while [245] found that the time delays between the hard and soft band variations may extend to very long periods (of the order of hundreds of days). A comparison between the graph on the right panel of Fig. 31 and the cross-correlation plot in the right panel of Fig. 30 reveals the power of studying the delays in the frequency domain instead of the time domain. The cross-correlation function indicates a constant delay between the hard- and soft-band variations, whereas the time lags reveal that the delays between the sinusoids that are responsible for the observed variations are not constant. Instead, the delays increase with increasing Fourier period, and for any given period, the time lag increases with the energy separation of the bands. As we mentioned, this behaviour can be explained by the Comptonization hypothesis, but in this case, we would expect that the delays between various energy bands should increase by the same amount for all Fourier periods, which is not what is observed [e.g. 17].
@@
This statistic is an estimate of the integral of the power spectrum from the shortest to the longest frequency sampled by the light curve. It is a biased estimator of this integral (even in the case of evenly sampled light curves), because of the red noise character of the intrinsic X-ray variations, and the bias depends on the PSD slope and the sampling pattern of the light curve. Its statistical properties have been studied by Vaughan et al. [307], in the case of evenly sampled light curves, and by [16], in all cases (also see Almaini et al. 11 and Buchner et al. 23 for alternative approaches in the low-count-rate regime). In general, a single measurement of the normalised excess variance is not very useful. Even when corrected for bias, single measurements follow a very asymmetric and broad distribution of unknown variance (that is, their error is unknown), especially in the case of sparse sampling and low signal-to-noise light curves. If we can calculate excess variance using light curves from many individual sources, then ’ensemble’ estimates of can be used, as they have a known error and distributions similar to a Gaussian [16]. The same procedure can be followed with excess variance measurements using segments of long light curves of a single object. However, the individual values in this case are not independent (due to the strong correlation between the points of the light curve over very long periods), therefore it will be difficult to determine the statistical properties of the resulting ensemble estimates.
@@
[236] computed for a large sample of (nearby) AGN using XMM-newton light curves, to study the X-ray variability of AGN in various energy bands and on timescales shorter than a day. They were able to show that their results could only be explained if the PSD bending frequency depends on both the BH mass and the accretion rate as shown by McHardy et al. [195, see the right panel of Fig. 32]. They also suggested that the PSD amplitude should increase with decreasing accretion rate in AGN. Although they predicted a rather steep dependence of the PSD amplitude on , which is not supported by other observations, their results showed that analysis can provide powerful insights into the properties of AGN X-ray variability. In fact, since can be used to characterise the X-ray variability of large samples of sources, the results of the studies can impose constraints on models for fundamental relations in AGN. For instance, [100] linked the demographics of AGN to their ensemble X-ray variability properties in order to investigate the black hole mass versus stellar mass relation in these objects (also see 282 for a similar application using PSD analysis).
@@
Given the tight correlation between and MBH, it was soon suggested that could be used to estimate the BH mass in AGN [e.g. 213, 209, 208]. [236] provided BH mass estimates for tens of AGN based on their measurements, and [334] provided a calibration for the correlation between and MBH in AGN. [14] critically assessed as a tool to measure MBH in AGN. They showed that it is possible to accurately measure MBH using measurements in the 3–10 and the 10–20 keV bands; however, the signal-to-noise ratio and the duration of the light curves should be larger than and ksec, respectively.
@@
All the works mentioned above used light curves that are shorter than day. Excess-variance measurements have perhaps played an even more important role in the study of the X-ray variability of high- AGN on longer timescales. In most cases, such objects are observed only a few times, in a highly irregular pattern. Consequently, the resulting light curves are not appropriate for the application of Fourier techniques. Therefore, has been widely used in the study of their X-ray variability properties. The lack of BH mass estimates was another obstacle in investigating the correlations between the amplitude of the X-ray variability and the physical parameters of AGN (i.e. MBH and ). As a result, most studies originally tried to investigate the correlation between variability amplitude and source luminosity, as well as redshift [e.g. 11, 172, 246, 222, 178, 322, 262, 321, 237, 332], confirming the anticorrelation with luminosity observed at lower redshift. This finding was also confirmed by studies based on SF analysis [312]. Almaini et al. [11], Manners et al. [172], Paolillo et al. [246] also observed a tentative evolution of X-ray variability with redshift, which was however disproved by later studies using better data and more complete samples.
@@ 3.9 Stationarity
It is well-known that the X-ray variability process in AGN is not strictly stationary because the variance of the lightcurves increases linearly with source flux. This “rms-flux” relation was first noticed by [301], and it is well established in AGN as well as in GBHBs [e.g. 118]. This is an important property and can place constraints on various models that have been proposed to explain the observed variations. For example, [305] showed that if the rms-flux relation holds for variations on all timescales, then the variability process in AGN should be multiplicative in nature: variations should be coupled together on all timescales. Therefore, the observed variability cannot be due to a shot-noise process or the result from completely independent variations in separate emitting regions.
@@ 3.10 Constrains on theoretical models
The model predicts that should increase with increasing energy because higher energy photons are produced at the hotter regions of the corona, which are closer to the BH, where is shorter. However, this does not appear to be the case with the energy-dependent PSDs in AGN (see Sect. 3.6.2). The bending frequency remains roughly the same throughout the observed frequency band from keV, which is against the predictions of the model. Furthermore, according to the model, the power spectra should extend with a slope of over a frequency range determined by up to 1/ (where the PSD break should appear). However, should be very large in order to explain power spectra such as the NGC 4051 PSD (left panel in Fig. 26). This power spectrum extends with a slope of over more than five orders in frequency down to Hz. See also the high-quality X-ray PSD of NGC 4395 in [33], which extends with a slope of below over four and a half orders of magnitudes in frequency. In order to explain this, the hot corona must extend to very large radii. For example, [253] fitted the Cyg X-1 PSD in its soft state with a fluctuations propagation model and found that the corona (which is located on top and below the disc) should have a very large radius of . This is clearly much larger than the size of the X-ray corona, as measured by monitoring observations of lensed quasars (see references at the beginning of Sect. 3). Apart from the difference between the model and the observed corona size, it is also difficult to understand how such a corona can be formed at such large radii from the BH, whether it will be stable or how it may evolve over time. Currently, there are no physical models that could be used to determine the accretion rate of the corona, its temperature, and the density radial distribution.
Note that the requirement of an extended corona does not come from the need to explain the observed long-term variations since these variations could be produced by accretion rate fluctuations which could propagate in the accretion disc and then within the corona as well). However, an accreting X-ray corona, with a large radius, is necessary to explain the long-term, continuum hard X-ray time-lags in AGN, especially if they really extend to very low frequencies [e.g. 245]. We also stress the similarity between the AGN PSDs and that of Cyg X-1 in its (own) soft state. In this state, a strong black-body-like component appears at low energies in the energy spectrum of the source. It is generally accepted that this component is due to thermal emission from the accretion disc, which extends down to the innermost stable circular orbit. If the similarity of the timing properties implies similar physical conditions as well, then it is possible that the accretion disc in the AGN also extends to its innermost radius, leaving not much space for a radially extended corona.